OpenStudy (anonymous):

Help Please, will give Medal! Factor the Trinomial below a^2 + 12ab + 27b^2

OpenStudy (anonymous):

\[a^2 + 12ab + 27b^2 =a^2 + 3ab +9ab+ 27b^2=3a(a+3b)+9b(a+3b)\] =(a+3b)(3a+9b) Hence (a+3b)(3a+9b) are the required factors of the given polynomial. @HeatDude

OpenStudy (anonymous):

okay thank you!!:)